OpenStudy (anonymous):

plz help me i am not good at this at all... i have 5 questions will medal

OpenStudy (anonymous):



Vocaloid (vocaloid):

are you sure you typed that equation correctly? there is no y in the equation

OpenStudy (anonymous):



OpenStudy (anonymous):

sry

Vocaloid (vocaloid):

ok, so our equation is 4x - 5y = 9 start by subtracting 4x from each side

OpenStudy (anonymous):

ok so -5y=5x

Vocaloid (vocaloid):

nope, not quite. remember that you cannot combine numbers and variables

OpenStudy (anonymous):

then i am not so sure

Vocaloid (vocaloid):

-5y = 9 - 4x the 9 and the 4x cannot be put together to get 5x, since 9 is a number and 5x is a variable

OpenStudy (anonymous):

ohhh.

Vocaloid (vocaloid):

now divide both sides by (-5)

OpenStudy (amistre64):

recall that 4x means: x+x+x+x 9-x-x-x-x is not the same as 5x, or x+x+x+x+x

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

@gahoolianowl

Vocaloid (vocaloid):

-5y = 9 - 4x dividing both sides by -5 gives us y = (9-4x)/(-5)